CARBON DIOXIDE CLUSTERS: (CO_{2})_{6} TO (CO_{2})_{13}

\underline{A.R.W. MCKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada; J. NOROOZ OLIAEE, M. DEHGHANY, and N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, AB T2N 1N4, Canada.

We recenty reported^{a} assignments of specific infrared bands in the CO_{2} ν_{3} region (∼2350 cm^{−1}) to (CO_{2})_{6}, (CO_{2})_{7}, (CO_{2})_{9}, (CO_{2})_{10}, (CO_{2})_{11}, (CO_{2})_{12}, and (CO_{2})_{13}. Spectra are obtained by direct absorption using a rapid-scan tuneable diode laser spectrometer to probe a pulsed supersonic slit-jet expansion and assignments are facilitated by recent calculations of Takeuchi based on the Murthy potential.^{b} (CO_{2})_{6} is a symmetric top with S_{6} point group symmetry which can be thought of as a stack of two planar cyclic trimers. (CO_{2})_{13} is also an S_{6} symmetric top, and consists of a single CO_{2} monomer surrounded by an slightly distorted icosahedral cage. The remaining clusters are asymmetric tops without symmetry. Here we report additional CO_{2} cluster results. Calculations based on the SAPT-s potential^{c} indicate that the structure of (CO_{2})_{10} may be slightly different from that given by Takeuchi/Murthy. An additional band is observed for each of (CO_{2})_{13} and (CO_{2})_{10}. A feature observed at 2378.2 cm^{−1} is assigned as a (CO_{2})_{6} parallel combination band involving the sum of a fundamental and a low-lying intermolecular vibration. Most significantly, two bands are assigned to a second isomer of (CO_{2})_{6}. This is also a symmetric top, but now with S_{4} symmetry. The two symmetric hexamer isomers observed spectroscopically correspond well with the lowest energy structures given by both the SAPT-s and Murthy intermolecular potentials.

aJ. Norooz Oliaee, M. Dehgany, N. Moazzen-Ahmadi, and A.R.W. McKellar, Phys. Chem. Chem. Phys. 13, 1297 (2011).

bH. Takeuchi, J. Phys. Chem. A 107, 5703 (2008); C.S. Murthy, S.F. O’Shea, and I.R. McDonald, Mol. Phys. 50, 531 (1983).

cR. Bukowski, J. Sadlej, B. Jeziorski, P. Jankowski, K. Szalewicz, S.A. Kucharski, H.L. Williams, and B.M. Rice, J. Chem. Phys.

110, 3785 (1999).